OpenStudy (anonymous):

Y = x - x^(-1) is a solution to which of the following differential equations? xy' + y = 2x xy' - y = 2x xy' + x = 2y xy' - x = 2y

OpenStudy (turingtest):

plug in for y and differentiate and see which holds

OpenStudy (anonymous):

would it be the first one?

OpenStudy (turingtest):

I don't know, why don't you show your work?

OpenStudy (anonymous):

the derivative of y is 1/x^2 +1 so xy' = (x^2+1)/x and (x^2+1)/x + x-(x^-1) = 2x

OpenStudy (turingtest):

yep, looks good

OpenStudy (anonymous):

thanks!

OpenStudy (turingtest):

welcome!